-/10 POINTS Find an equation of the tangent line for the curve x=tet, y=t+et at the point corresponding to t=0.
Answer #1

Given, x=tet i y=ttet ..:** (t)=tet . + y(t) = t tet At t=0,4 (0)=0.e=0 At t=0, y10) = 0 +e=1 Now, do te 1 yz tret (tet) olen